# Predicting Energy Demand with Tensor Factor Models

### v1

@@ Abstract
European Central Bank Università di Bologna Università di Bologna
This version: February 10, 2025111Disclaimer: the views expressed in this paper are those of the authors and do not necessarily reflect the views and policies of the European Central Bank or the Eurosystem.
@@ 2 Electricity demand: the PJM dataset
The data used in this work are produced by PJM Interconnection LLC. PJM is a Regional Transmission Organization (RTO) coordinating energy distribution across 13 U.S. states and the District of Columbia. [^2] More information available at https://pjm.com/. For geographical reference, the 13 States are all located between the eastern part of the Midwest, the northern part of the Southeast, and the southern part of the Northeast. [^3] The States totally or partially covered by PJM are Delaware, Illinois, Indiana, Kentucky, Maryland, Michigan, New Jersey, North Carolina, Ohio, Pennsylvania, Tennessee, Virginia, West Virginia, plus the District of Columbia. The raw datasets used in this work are freely available at this link. [^4] https://kaggle.com/datasets/robikscube/hourly-energy-consumption/data In total, data are available for 11 different electricity providers or groups of providers. [^5] The geographical location of the providers is available under the Transmission Zones button at this link and are labeled in the following manner: American Electric Power (AEP), Commonwealth Edison (ComEd), Dayton Power and Light Company (DAYTON), Duke Energy Ohio/Kentucky (DEOK), Dominion Virginia Power (DOM), Duquesne Light Co. (DUQ), East Kentucky Power Cooperative (EKPC), FirstEnergy (FE), Northern Illinois Hub (NI), PJM East Region (PJME), PJM East Region (PJMW). In the latter two cases, PJME and PJMW are not unique providers but aggregates. Hourly data between 2002 and 2018 in Megawatts (MW) are included. As most of the series does not record measurements starting from 2002, we only retain information on nine proiders (AEP, ComEd, DAYTON, DEOK, DOM, DUQ, FE, PJME, PJMW) from the beginning of July 2012 to the end of July 2018, collecting 57,456 hourly observations for each provider. Figure 1 displays the time series of the different providers. Despite data points being very dense, we can highlight a few interesting aspects that are relevant for model definition. Firstly, levels differ greatly between datasets. This is consistent with the geographical area of energy supply for the different providers: some companies operate in much larger areas than others. Secondly, the behavior of the series is strongly seasonal, and the absence of long-run trends is well-explained by the relatively short time period covered by the series (about 7 years). Some peaks are evident. Peaks are similar in magnitude and their distance from the others is regular, within the same time series. Across time series, the peaks are typically aligned over time. Interestingly, time series such as AEP, DEOK, and DOM present extremely regular over-the-mean peaks (regular-peaked series). On the contrary, other time series such as COMED and DUQ present alternated over-the-mean peaks with different magnitude levels (alternate-peaked series). All 9 time series can be easily categorized into one of these two categories. A closest qualitative analysis of the time series shows that all peaks (both for regular-peaked series and alternate-peak series) are semiannual. This is consistent with the typical behavior of electricity consumption. Indeed, semestral peaks correspond to the hottest weeks in summer, and the coldest weeks in winter. Typically, for the alternated peak series, the highest peaks are the ones in summer. This is probably due to the wide adoption of air conditioning in the U.S.A. for cooling offices, houses, and factories. On the other hand, heating systems do not only consume electricity but also natural gas. In general, even in regular-peaked time series summer peaks are slightly higher than winter peaks. This is further confirmed by the seasonal plot displayed in Figure 2 showing the weakly average electricity demand for the companies, normalized by total company demand.
@@ 3.1 A General Modeling Framework
$\begin{array}[]{rll}y_{i,\mathbf{s},t}&=&\mu_{i,\mathbf{s}}+\sigma_{i,\mathbf{s}}\epsilon_{i,\mathbf{s},t},\\ \epsilon_{i,\mathbf{s},t}&=&\sum\limits_{r=1}^{R}{\lambda}_{i,r}{f}^{(M)}_{r,\mathbf{s},t}+\nu_{i,\mathbf{s},t},\\ f^{(M)}_{r,\mathbf{s},t}&=&\sum\limits_{k_{M}=1}^{K_{M}}{\beta}^{(M)}_{s_{M},k_{M}}{f}^{(M-1)}_{r,s_{1},\dots,k_{M},t}+\eta^{(M)}_{r,\mathbf{s},t},\\ {f}^{(M-1)}_{r,s_{1},\dots,s_{M-1},k_{M},t}&=&\sum\limits_{k_{M-1}=1}^{K_{M-1}}{\beta}^{(M-1)}_{s_{M-1},k_{M-1}}f^{(M-2)}_{r,s_{1},\dots,k_{M-1},k_{M},t}+\eta^{(M-1)}_{r,s_{1},\dots,s_{M-1},k_{M},t},\\ &\vdots&\\ f^{(1)}_{r,s_{1},k_{2}\dots,k_{M},t}&=&\sum\limits_{k_{1}=1}^{K_{1}}{\beta}^{(1)}_{s_{1},k_{1}}{f}_{r,k_{1},\dots,k_{M},t}+\eta^{(1)}_{r,s_{1},k_{2}\dots,k_{M},t}.\end{array}$ (1)
@@
$\begin{array}[]{rll}y_{i,s_{1},s_{2},t}&=&\mu_{i,s_{1},s_{2}}+\sigma_{i,s_{1},s_{2}}\epsilon_{i,s_{1},s_{2},t}\\ \epsilon_{i,s_{1},s_{2},t}&=&{\lambda}_{i}{f}^{(2)}_{s_{1},s_{2},t}+\nu_{i,s_{1},s_{2},t},\\ f^{(2)}_{s_{1},s_{2},t}&=&{\beta}^{(2)}_{s_{2}}{f}^{(1)}_{s_{1},t}+\eta^{(2)}_{s_{1},s_{2},t},\\ f^{(1)}_{s_{1},t}&=&{\beta}^{(1)}_{s_{1}}{f}_{t}+\eta^{(1)}_{s_{1},t},\end{array}$
@@
$\begin{array}[]{rll}y_{i,s_{1},s_{2},t}&=&\mu_{i,s_{1},s_{2}}+\sigma_{i,s_{1},s_{2}}\epsilon_{i,s_{1},s_{2},t}\\ \epsilon_{i,s_{1},s_{2},t}&=&{\lambda}_{i}{\beta}^{(2)}_{s_{2}}{\beta}^{(1)}_{s_{1}}{f}_{t}+\varepsilon_{i,s_{1},s_{2},t},\\ \varepsilon_{i,s_{1},s_{2},t}&=&{\lambda}_{i}\left({\beta}^{(2)}_{s_{2}}\eta^{(1)}_{s_{1},t}+\eta^{(2)}_{s_{1},s_{2},t}\right)+\nu_{i,s_{1},s_{2},t}.\\ \end{array}$
@@ 3.3 Estimation
~$\widetilde{\mathcal{F}}_{t}=\frac{1}{NS}\,\mathcal{X}_{t}\times^{M}_{j=1}\widetilde{\mathbf{B}}^{(j)}\times_{M+1}\widetilde{\bm{\Lambda}}.$
@@
where ~$\widetilde{\mathcal{M}}=[\widehat{\mu}_{i,\mathbf{s}}]$, $\widetilde{\mathcal{S}}=[\widehat{\sigma}_{i,\mathbf{s}}]$ and $\odot$ denotes the element-wise (Hadamard) product.
@@
Output: Factor loading matrices $\widetilde{\bm{\Lambda}}$ and $\{\widetilde{\mathbf{B}}^{(j)},1\leq j\leq M\}$
@@
$\begin{array}[]{rll}\widehat{\mathbf{B}}&=&\sqrt{S}\times\text{the $R$ eigenvectors of }\frac{1}{TNS}\sum\limits_{t=1}^{T}\mathbf{X}^{\prime}_{1,t}\mathbf{X}_{1,t},\\[7.22743pt] \widehat{\bm{\Gamma}}^{(j)}&=&\sqrt{NS_{-j}}\times\text{the $(RK_{-j})$ eigenvectors of }\frac{1}{TNS}\sum\limits_{t=1}^{T}\mathbf{X}^{\prime}_{j+1,t}\mathbf{X}_{j+1,t},\quad 1\leq j\leq M.\\[7.22743pt] \end{array}$
@@
$\begin{array}[]{rll}\widetilde{\bm{\Lambda}}&=&\sqrt{N}\times\text{the $R$ eigenvectors of }\frac{1}{TNS}\sum\limits_{t=1}^{T}\mathbf{X}_{1,t}\frac{\widehat{\mathbf{B}}^{\prime}\widehat{\mathbf{B}}}{S}\mathbf{X}^{\prime}_{1,t},\\[7.22743pt] \widetilde{\mathbf{B}}^{(j)}&=&\sqrt{S_{j}}\times\text{the $(K_{j})$ eigenvectors of }\frac{1}{TNS}\sum\limits_{t=1}^{T}\mathbf{X}_{j+1,t}\frac{\widehat{\bm{\Gamma}}^{(j)\prime}\widehat{\bm{\Gamma}}^{(j)}}{S_{-j}}\mathbf{X}^{\prime}_{j+1,t},\quad 1\leq j\leq M.\\[7.22743pt] \end{array}$
@@
[…]
@@ 3.4 Forecasting
Once the factor dynamics have been estimated, future values of the factor tensor, denoted as ~$\widetilde{\mathcal{F}}_{T+n|T}$, are obtained through recursive forecasting. Given the predicted $n$-step-ahead factor tensor, forecasts of the original observed data $\mathcal{Y}_{t}$ are obtained by plugging ~$\widetilde{\mathcal{F}}_{T+n|T}$ into the estimated factor model structure. Specifically, the $n$-step-ahead forecast is computed as
@@ 4.2 Forecast evaluation
where $w$ is the window of reference, $n$ is the horizon of the forecast, $W=T-n$ is the total number of rolling windows and $\widetilde{\mathbf{Y}}_{i,t+n|t}$ is the $n$-step ahead forecast.

### v2

@@ Abstract
Anima Sgr Università di Bologna Università di Bologna
This version: March, 2026
00footnotetext: Disclaimer: the views expressed in this paper are those of the authors and do not necessarily reflect the views and policies of Anima Sgr. 00footnotetext: Acknowledgments: LT acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 1409 published on 14.9.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union –NextGenerationEU– Project Title “Measuring, managing and hedging indirect climate-transition risks” –CUP P20228CHNL– Grant Assignment Decree No. 1376 adopted on 01/09/2023 by the Italian Ministry of Ministry of University and Research (MUR).
@@ 2 Electricity demand: the PJM dataset
The data used in this work are produced by PJM Interconnection LLC. PJM is a Regional Transmission Organization (RTO) coordinating energy distribution across 13 U.S. states and the District of Columbia. [^1] More information available at https://pjm.com/. For geographical reference, the 13 States are all located between the eastern part of the Midwest, the northern part of the Southeast, and the southern part of the Northeast. [^2] The States totally or partially covered by PJM are Delaware, Illinois, Indiana, Kentucky, Maryland, Michigan, New Jersey, North Carolina, Ohio, Pennsylvania, Tennessee, Virginia, West Virginia, plus the District of Columbia. The raw datasets used in this work are freely available at this link. [^3] https://kaggle.com/datasets/robikscube/hourly-energy-consumption/data In total, data are available for 11 different electricity providers or groups of providers. [^4] The geographical location of the providers is available under the Transmission Zones button at this link and are labeled in the following manner: American Electric Power (AEP), Commonwealth Edison (ComEd), Dayton Power and Light Company (DAYTON), Duke Energy Ohio/Kentucky (DEOK), Dominion Virginia Power (DOM), Duquesne Light Co. (DUQ), East Kentucky Power Cooperative (EKPC), FirstEnergy (FE), Northern Illinois Hub (NI), PJM East Region (PJME), PJM East Region (PJMW). In the latter two cases, PJME and PJMW are not unique providers but aggregates. Hourly data between 2002 and 2018 in Megawatts (MW) are included. As most of the series does not record measurements starting from 2002, we only retain information on nine proiders (AEP, ComEd, DAYTON, DEOK, DOM, DUQ, FE, PJME, PJMW) from the beginning of January 2012 to the end of July 2018, collecting 57,456 hourly observations for each provider. Figure 1 displays the time series of the different providers. Despite data points being very dense, we can highlight a few interesting aspects that are relevant for model definition. Firstly, levels differ greatly between datasets. This is consistent with the geographical area of energy supply for the different providers: some companies operate in much larger areas than others. Secondly, the behavior of the series is strongly seasonal, and the absence of long-run trends is well-explained by the relatively short time period covered by the series (about 7 years). Some peaks are evident. Peaks are similar in magnitude and their distance from the others is regular, within the same time series. Across time series, the peaks are typically aligned over time. Interestingly, time series such as AEP, DEOK, and DOM present extremely regular over-the-mean peaks (regular-peaked series). On the contrary, other time series such as COMED and DUQ present alternated over-the-mean peaks with different magnitude levels (alternate-peaked series). All 9 time series can be easily categorized into one of these two categories. A closest qualitative analysis of the time series shows that all peaks (both for regular-peaked series and alternate-peak series) are semiannual. This is consistent with the typical behavior of electricity consumption. Indeed, semestral peaks correspond to the hottest weeks in summer, and the coldest weeks in winter. Typically, for the alternated peak series, the highest peaks are the ones in summer. This is probably due to the wide adoption of air conditioning in the U.S.A. for cooling offices, houses, and factories. On the other hand, heating systems do not only consume electricity but also natural gas. In general, even in regular-peaked time series summer peaks are slightly higher than winter peaks. This is further confirmed by the seasonal plot displayed in Figure 2 showing the weakly average electricity demand for the companies, normalized by total company demand.
@@ 3.1 A General Modeling Framework
[…]
@@
[…]
@@
[…]
@@ 3.3 Estimation
$\widetilde{\mathcal{F}}_{t}=\frac{1}{NS}\,\mathcal{X}_{t}\times^{M}_{j=1}\widetilde{\mathbf{B}}^{(j)}\times_{M+1}\widetilde{\bm{\Lambda}}$.
@@
where $\widetilde{\mathcal{M}}=[\widehat{\mu}_{i,\mathbf{s}}]$, $\widetilde{\mathcal{S}}=[\widehat{\sigma}_{i,\mathbf{s}}]$ and $\odot$ denotes the element-wise (Hadamard) product.
@@
1
0.5em Output: Factor loading matrices $\widetilde{\bm{\Lambda}}$ and $\{\widetilde{\mathbf{B}}^{(j)},1\leq j\leq M\}$
@@
[…]
@@
[…]
@@
Note that if the process in (4) satisfies Assumptions 1-3 in Yu et al. (2022), there exists matrices $\widehat{\mathbf{H}}$ and $\widehat{\mathbf{J}}_{j}$ satisfying $\widehat{\mathbf{H}}^{\prime}\widehat{\mathbf{H}}\xrightarrow{p}\mathbb{I}_{R}$ and $\widehat{\mathbf{J}}^{\prime}_{j}\widehat{\mathbf{J}}_{j}\xrightarrow{p}\mathbb{I}_{K_{j}}$, with $j=1,\dots,M$, such that,
@@ 3.4 Forecasting
Once the factor dynamics have been estimated, future values of the factor tensor, denoted as $\widetilde{\mathcal{F}}_{T+n|T}$, are obtained through recursive forecasting. Given the predicted $n$-step-ahead factor tensor, forecasts of the original observed data $\mathcal{Y}_{t}$ are obtained by plugging $\widetilde{\mathcal{F}}_{T+n|T}$ into the estimated factor model structure. Specifically, the $n$-step-ahead forecast is computed as
@@ 4.2 Forecast evaluation
where $w$ is the window of reference, $n$ is the horizon of the forecast, W=TnW=T-n is the total number of rolling windows and $\widetilde{\mathbf{Y}}_{i,t+n|t}$ is the $n$-step ahead forecast.